please help me.
A ball is thrown upwards from the roof of a buildiung 100 m tall with an inituial velocity of 20m/s. When will the ball reach a height of 80m?
I believe the equation is h=5t^2+80t+100. How do I solve using quadratic formula?
Thanks

Answer by Nate(3500)

f(x) = -16t^2 + 20t + 100
~> Ball reaches 80m.
80 = -16t^2 + 20t + 100
-20 = -16t^2 + 20t
5/4 = t^2 - 5t/4
105/4 = 80/64 + 25/64 = (t - 5/8)^2
+-sqrt(105)/2 = t - 5/8
5/8 +- 4*sqrt(105)/8 = t